OpenStudy (anonymous):

Factor 25x^2 + 10x + 1. (25x + 1)(x + 1) (5x + 1)(5x - 1) (5x + 1)²

OpenStudy (anonymous):

so if i cant combine like terms then what do i do

OpenStudy (campbell_st):

this expression is a quadratic.... always in the form \[ax^2 + bx + c\] the easy way to solve these is as follows next write your expression is n brackets with \[\frac{(ax + factor1)(ax+factor2)}{a}\] \[\frac{(25x+5)(25x+5)}{25}\] look for common factors in both pairs of brackets \[\frac{5\times5\times(5x+1)(5x+1)}{25}\] cancel the common factor of 25 leaving \[(5x +1)(5x+1) = (5x+1)^2\]